Hello my name is Sarah I was here yesterday and you guys did me a big favor!! but now all I have is 5 more problems. {y=5x 6x-2y=-4 {4x-y=5 x=8-2y {-7x+y=32 2x+3y=27 {2x+y=5 4x-3y=-10 {4x+2y=7 x-5y=10 can i have it y tonight? thanks (:
Found 2 solutions by ewatrrr, lynnlo:
Answer by ewatrrr(24785)

 
Hi,
Important that you learn how to do these:
y=5x
6x-2y=-4 ||Substitute for y and find x ( then y = 5x)
x=8-2y+ ||Substitute for x and find y (then x = 8-2y)
4x-y=5
-7x+y=32 || Multiply 1st EQ by -3 and add to eliminate y (you'll find x =-3
2x+3y=27
2x+y=5 || Multiply 1st EQ by 3 and add to eliminate y (you'll find x =1/2
4x-3y=-10
4x+2y=7
x-5y=10 || Multiply 2nd EQ by -4 and add to eliminate x (you'll find y = -3/2

Answer by lynnlo(4176)

(1)x=1,y=5
(2)x=8/3,y=-8/3
(3)x=-3,y=11
(4)x=1/2,y=-3/2
(5)x=5/2,y=-3/2